# Remote sensing colour image semantic segmentation of trails created by large herbivorous mammals

### v1

@@ Abstract
Detection of spatial areas where biodiversity is at risk is of paramount importance for the conservation and monitoring of ecosystems. Large terrestrial mammalian herbivores are keystone species as their activity not only has deep effects on soils, plants, and animals, but also shapes landscapes, as large herbivores act as allogenic ecosystem engineers. One key landscape feature that indicates intense herbivore activity and potentially impacts biodiversity is the formation of grazing trails. Grazing trails are formed by the continuous trampling activity of large herbivores that can produce complex networks of tracks of bare soil. Here, we evaluated different algorithms based on machine learning techniques to identify grazing trails. Our goal is to automatically detect potential areas with intense herbivory activity, which might be beneficial for conservation and management plans.
We have applied five semantic segmentation methods combined with fourteen encoders aimed at mapping grazing trails on aerial images. Our results indicate that in most cases the chosen methodology successfully mapped the trails, although there were a few instances where the actual trail structure was underestimated. The UNet architecture with the MambaOut encoder was the best architecture for mapping trails. The proposed approach could be applied to develop tools for mapping and monitoring temporal changes in these landscape structures to support habitat conservation and land management programs. This is the first time, to the best of our knowledge, that competitive image segmentation results are obtained for the detection and delineation of trails of large herbivorous mammals.

### v2

@@ Abstract
Identifying regions where biodiversity faces elevated risk is essential for guiding conservation efforts and ecosystem monitoring. Large terrestrial herbivores act as allogenic engineers: through their feeding and trampling, they reshape soils, vegetation, and animal communities. One of the most visible signs of intense herbivore presence is the formation of grazing trails—interconnected bare-soil pathways created by repeated trampling. In this work, we benchmarked five semantic segmentation architectures, each combined with fourteen different encoder backbones, to automatically pinpoint grazing trails in aerial photographs. Our goal was to highlight areas of concentrated herbivory, thereby providing actionable information for habitat management and conservation planning. Most model-encoder pairings successfully delineated the trail networks, although a few combinations slightly underestimated their full extent. Overall, the UNet framework paired with the MambaOut encoder achieved the highest accuracy in mapping these paths. The pipeline we introduce can be integrated into tools that track the emergence and evolution of grazing trails over time, offering valuable support for land-use management and biodiversity protection programs. To our knowledge, this represents the first instance of achieving competitive image segmentation performance for the detection and precise outlining of large herbivore trail systems.